### v2

# On Bureau’s classification of quadratic differential equations in two variables free of movable critical points

Abstract.
As part of the efforts aimed at extending Painlevé and Gambier’s work on second-order equations in one variable to first-order ones in two, in 1981, Bureau classified the systems of ordinary quadratic differential equations in two variables which are free of movable critical points (which have the Painlevé Property). We revisit this classification, which we complete by adding some cases overlooked by Bureau, and by correcting some of his arguments. We also simplify the canonical forms of some systems, bring the natural symmetries of others into their study, and investigate the birational equivalence among some of the systems in the class. Lastly, we study the birational geometry of Okamoto’s space of initial conditions for Bureau’s system VIII, in order to establish the sufficiency of some necessary conditions for the absence of movable critical points.
Key words and phrases:
movable critical points, Painlevé Property, Painlevé transcendent, quadratic equation
2020 Mathematics Subject Classification: 34M04, 34M55
1. Introduction
Within the study of ordinary differential equations in the complex domain, the following problem is attributed to Goursat: to classify the systems of equations
$\begin{split}y^{\prime}&=P(y,z,t),\\ z^{\prime}&=Q(y,z,t),\end{split}$ (1)
with $P$ and $Q$ rational in $y$ and $z$, and analytic in the independent variable $t$, which are free of movable critical points, with two systems differing by a transformation which is analytic in the independent variable and birational in the dependent ones considered equal [Gar60]. Following Painlevé, a critical point of a function is a singular point around which at least two of its branches are permuted (see the first of his Leçons de Stockholm [Pai97]). The above is a natural and broad generalization of the problem solved by Painlevé and Gambier concerning second-order equations in one variable, which lead to the discovery of the Painlevé transcendents. Garnier wondered if such general systems could define transcendents beyond those that appear in Painlevé and Gambier’s classification, and estimated this to be a difficult problem. One may restrict it by considering systems (1) with $P$ and $Q$ polynomial in $y$ and $z$. These polynomial equations already form a rich class, which naturally contains Painlevé’s first two, and which, as shown by Malmquist [Mal23, §33], contains the other Painlevé equations as well (see also [Oka79]). Goffar-Lombet investigated the question for polynomial systems having some particular homogeneous polynomials, of degrees four and five, as highest-degree terms, obtaining precise and complete results for such systems [GL74]. In 1980, Kimura and Matuda announced a classification of polynomial systems of the form (1) free of movable critical points, of degree up to five [KM80], which does not seem to have had a follow-up from their part. Shortly afterwards, in a series of articles published in the Bulletin de la Classe des Sciences of the Académie Royale de Belgique ([Bur81a], [Bur81b], [Bur81c]), Bureau classified quadratic systems up to affine transformations. His results are summarized in Table III in [Bur81c, §28] (he uses the term stable to refer to systems that are free of movable critical points). He later studied higher degree polynomial systems in the subsequent articles of the series. The aim of this article is to revisit and revise Bureau’s classification of the quadratic case:
Theorem 1.
Consider a system of differential equations of the form (1), with $P$ and $Q$ polynomials in $y$ and $z$ of degree at most two (with at least one of them non-linear), whose coefficients are holomorphic functions defined in a domain $U\subset\mathbf{C}$. If the system is free of movable critical points then, for every $t_{0}$ in $U$, with the possible exception of a discrete set, there exists a neighborhood $V$ of $t_{0}$ within $U$, and a transformation $\phi:\mathbf{C}^{2}\times V\to\mathbf{C}^{2}\times\mathbf{C}$,
$(\zeta,t)\mapsto(\xi(\zeta,t),\tau(t))$
with $\xi$ affine in $\zeta$ and holomorphic in $t$, and $\tau$ a biholomorphism onto its image, such that the system solved by $\xi$, as a function of $\tau$, is one of those appearing in Table 1.
Reciprocally, all systems in Table 1 are free of movable critical points. They are integrated either in an elementary way or via some Poincaré transcendent (also indicated in Table 1).
System Equations Integration
I
$y^{\prime}=-y^{2}+az$
$z^{\prime}=-yz$
II
$y^{\prime}=y(y-2z)+a$
$z^{\prime}=-z^{2}$
III
$y^{\prime}=y^{2}$
$z^{\prime}=2yz+ay$
IV
$y^{\prime}=-y^{2}+az$
$z^{\prime}=0$
$\text{V}_{0}$
$y^{\prime}=0$
$z^{\prime}=6y^{2}$
V
$y^{\prime}=z$
$z^{\prime}=6y^{2}+f$
   with $f^{\prime\prime}=0$
Painlevé I
VI
$y^{\prime}=a$
$z^{\prime}=z(z+y)+b$
$\text{VII}_{0}$
$y^{\prime}=0$
$z^{\prime}=yz+a$
VII
$y^{\prime}=z$
$z^{\prime}=yz+a$
$\text{VIII}(n)$
($n\geq 1$)
$y^{\prime}=y[y-(n+2)z]+a$
$z^{\prime}=-z(ny+z)+b$
   with a condition of order $n-1$ (Section 3.6)
[…]
($n>1$)
$y^{\prime}=-y^{2}$
$z^{\prime}=-(n+1)yz+ay$
[…]
$y^{\prime}=-y^{2}+z$
$z^{\prime}=-4yz+a$
[…]
($n>1$)
$y^{\prime}=-y^{2}$
$z^{\prime}=(n-1)yz+py$
   with $\mathrm{d}^{n-1}p/\mathrm{d}t^{n-1}=0$
[…]
$y^{\prime}=-y^{2}+z+12q$
$z^{\prime}=yz$
   with $(q^{\prime\prime}-6q^{2})^{\prime\prime}=0$
Painlevé I
[…]
$y^{\prime}=-y^{2}+z-\frac{1}{2}f$
$z^{\prime}=2yz+H$
   with $f^{\prime\prime}=0$
Painlevé II
[…]
$y^{\prime}=-y^{2}+z+3q$
$z^{\prime}=4yz-9q^{2}$
   with $(q^{\prime\prime}-6q^{2})^{\prime\prime}=0$
Painlevé I
XI
$y^{\prime}=y(y-z)+ay$
$z^{\prime}=z(z-y)-az$
Table 1. Representatives of the quadratic systems free of movable critical points (Theorem 1). Here, $n\in\mathbf{Z}$, $a$, $b$, etc. are functions, $H$ and $K$ are constants.
System Equations Int.
XII
$y^{\prime}=y(2z+y)+2fy-H$
$z^{\prime}=-z(2y+z)-2fz+K$
   with $f^{\prime\prime}=0$
P. IV
XIII
$y^{\prime}=\frac{1}{2}y(2z-y)+2py$
$z^{\prime}=\frac{1}{2}z(3y-2z)-4pz+2p^{2}-2p^{\prime}+f$
   with $f^{\prime\prime}=0$, $(p^{\prime\prime}-p^{3}-fp)^{\prime}=0$
P. II
XIV
$y^{\prime}=y(2z-y)+3py+r$
$z^{\prime}=z(y-z)-2pz$
   with $r^{\prime}=pr$, and, for $q=\frac{1}{12}(p^{\prime}+p^{2}-r)$, $(q^{\prime\prime}-6q^{2})^{\prime\prime}=0$
P. I
Table 1 (continued).
[…]
Some systems belonging to different rows of Table 1 are birationally equivalent, or are birationally equivalent to some Painlevé equation. The following result will be proved in the course of establishing Theorem 1.
Proposition 2.
We have the following birational equivalences.
  1. (1)
    A system
    • IX.B(2) with parameter $q$; or
    • IX.B(5) with parameter $q$; or
    • XIV with parameters $(r,p)$, for $q=\frac{1}{12}(p^{\prime}+p^{2}-r)$;
    is equivalent to the system V with parameter $f=q^{\prime\prime}-6q^{2}$ by a bimeromorphic map $\phi:\mathbf{C}^{2}\times\mathbf{C}\dashrightarrow\mathbf{C}^{2}\times\mathbf{C}$ of the form $(\zeta,t)\mapsto(\xi(\zeta,t),t)$, with $\xi$ birational as a function of $\zeta$ for most values of $t$. (System V is the first-order system associated to the Painlevé I equation and some of its degenerations.)
  2. (2)
    A system
    • XIII, with parameters $(f,p)$, for $\alpha=2p^{3}+fp-p^{\prime\prime}$; or
    • IX.B(3), with parameters $(f,H)$, for $\alpha=H-\frac{1}{2}f^{\prime}$;
    is birationally equivalent, by a bimeromorphic map as above, to the cubic first-order system associated to the equation
    $y^{\prime\prime}=2y^{3}+fy+\alpha\quad(f^{\prime\prime}=0,\;\alpha\in\mathbf{C}),$
    giving the Painlevé II equation and some of its degenerations.
  3. (3)
    A system XII, with parameters $(f,H,K)$, for $\alpha=\frac{1}{2}H-K-f^{\prime}$ and $\beta=-H^{2}$, is birationally equivalent, by a bimeromorphic map as above, to the rational first-order system associated to the equation
    $y^{\prime\prime}=\frac{(y^{\prime})^{2}}{2y}+\frac{3}{2}y^{3}+4fy^{2}+2(f^{2}-\alpha)y-\frac{\beta^{2}}{2y}\quad(f^{\prime\prime}=0,\;\alpha,\beta\in\mathbf{C}),$ (2)
    giving the Painlevé IV equation and some of its degenerations.
This implies that these systems are more than just free of movable critical points:
Corollary 3.
The solutions to the systems V, IX.B(2), IX.B(3), X.B(5), XII, XIII and XIV of Table 1 are single-valued meromorphic functions defined in all of $\mathbf{C}$.
The less satisfactory part of the classification concerns system VIII, the only one for which the necessary and sufficient conditions for the absence of movable critical points do not explicitly appear, neither in Bureau’s work, nor in Table 1. These systems form an infinite family, indexed by a positive integer $n$, and the complexity of a certain necessary condition for the absence of movable critical points grows with $n$; it seems difficult to give an explicit expression for this condition for every $n$. At the same time, the equations are, conceptually, very simple, as they factor through Riccati equations. In any case, there remains the question of whether the above necessary condition is also a sufficient one. We will show that this is indeed the case, by studying the birational geometry of these systems, in the spirit of Okamoto’s work on the spaces of initial conditions of the Painlevé equations  [Oka79], which is, on its turn, motivated by Painlevé’s 17th of his Leçons de Stockholm [Pai97]. This will be done in Section 3.6.
Remark 4.
The equations in Table 1 are not in their most reduced form. For a given system, there may be different parameter sets giving equivalent equations. We have favored the simplicity of the forms, at the cost of some minor redundancies.
Remark 5.
There is no system X in Table 1. Table III in [Bur81c] contains a “system X,” which summarizes the results in [Bur81b, §15]. There, under a different normalization, the analysis in [Bur81b, §14], the one leading to system VI, is duplicated. Bureau uses this to establish the connections with other equations that he had previously studied, but this does not contribute towards a classification.
This article can be read independently of Bureau’s work, although we have included information aimed at facilitating the parallel reading of both.
2. Criteria
We now describe some of the tools used by Bureau to establish the absence of movable critical points for the equations under consideration.
2.1. Quadratic homogeneous equations and univalence
Consider a system of equations of the form (1),
$\zeta^{\prime}=\sum_{i=0}^{n}P_{i}(\zeta,t),$
with $\zeta\in\mathbf{C}^{2}$ and $P_{i}$ a holomorphic function which, as a function of $\zeta$, is a homogeneous polynomial of degree $i$. For $\alpha\neq 0$, $\xi=\alpha z$, and $t=t_{0}+\alpha^{n-1}\tau$,
$\frac{\mathrm{d}\xi}{\mathrm{d}\tau}=\sum_{i=0}^{n}\alpha^{n-i}P_{i}(\xi,t_{0}+\alpha\tau).$
The limit as $\alpha\to 0$ is the autonomous, homogeneous equation $\mathrm{d}\xi/\mathrm{d}\tau=P_{n}(\xi,t_{0})$. If the original equation is free of movable critical points, the latter will have only single-valued solutions [Pai00] (see also [Bur80]). Thus, the first problem, in our setting, is to determine the equations of the form
$\begin{split}y^{\prime}&=P(y,z),\\ z^{\prime}&=Q(y,z),\end{split}$ (3)
with $P$ and $Q$ quadratic homogeneous polynomials, having single-valued solutions. This has been done by Goffar-Lombet in [GL74, Ch. I], by Bureau in [Bur81a], and by Ghys and Rebelo in [GR97, Thm. C] and [GR00] (and is likely to have been done elsewhere as well). The classification appears in Table 2, where we have listed representatives for all the linear equivalence classes of univalent quadratic homogeneous equations of the form (3); we follow Bureau’s normalizations. Observe that these representatives have no continuous parameters in them.
System reduced equation indices
I $\begin{array}[]{l}y^{\prime}=-y^{2}\\ z^{\prime}=-yz\end{array}$
V $\begin{array}[]{l}y^{\prime}=0\\ z^{\prime}=6y^{2}\end{array}$
III ($n=-1$)
IV ($n=1$)
IX ($n\notin\{-1,0,1\}$)
$\begin{array}[]{l}y^{\prime}=-y^{2}\\ z^{\prime}=(n-1)yz\end{array}$ $n$
VII $\begin{array}[]{l}y^{\prime}=0\\ z^{\prime}=yz\end{array}$infty
II ($n=0$)
VIII ($n>0$)
$\begin{array}[]{l}y^{\prime}=y[y-(n+2)z]\\ z^{\prime}=-z(ny+z)\end{array}$ $11,n+1,-n-$ 1
VI $\begin{array}[]{l}y^{\prime}=0\\ z^{\prime}=z(y+z)\end{array}$ $1,\infty,\infty$
XI $\begin{array}[]{l}y^{\prime}=y(y-z)\\ z^{\prime}=z(z-y)\end{array}$ $2,2,\infty$
XII $\begin{array}[]{l}y^{\prime}=y(2z+y)\\ z^{\prime}=-z(2y+z)\end{array}$ $3,3,3$
XIII $\begin{array}[]{l}y^{\prime}=\frac{1}{2}y(2z-y)\\ z^{\prime}=\frac{1}{2}z(3y-2z)\end{array}$ $2,4,4$
XIV $\begin{array}[]{l}y^{\prime}=y(2z-y)\\ z^{\prime}=z(y-z)\end{array}$ $2,3,6$
Table 2. Representatives of the univalent quadratic homogeneous systems (Bureau’s normalization), with the indices of their simple radial orbits
An approach to this classification is the following. For a generic solution $(y,z)$ of (3), $s=y/z$ is either constant, or satisfies an equation of the form
$-\frac{s^{\prime\prime}}{(s^{\prime})^{2}}=F(s),$ (4)
for some rational function $F$ (a similar equation is satisfied by $1/s$). Briot and Bouquet gave necessary and sufficient conditions on $F$ for the univalence of a function satisfying such an equation [BB55]: at the finite values of $s$, $F$ has only simple poles, and their residues are either equal to $-1$, or are of the form $11/k-$ 1 for $k\in\mathbf{Z}$, plus the analogous condition at $s=\infty$ coming from the equation satisfied by $1/s$ (see also Painlevé [Pai00, §10]; see [Gui06, Section 2.2] for an interpretation in terms of affine structures). The index of the pole of (4) is the associated integer $k$ (we will say that a simple pole with residue $-1$ has index $\infty$). We refer the reader to the closing remarks in [Con99, §2.3] for a discussion around this terminology.
A radial orbit for an equation of the form (3) is a line through $(0,0)$ that is invariant by the equation (that is a union of solutions). If not every line through the origin is a radial orbit, the equation has three of them (counted with multiplicity). The poles of the associated equation (4) come from radial orbits, although some radial orbits may fail to produce them. The index of a radial orbit is the one of the corresponding pole in equation (4); a simple radial orbit which produces no poles for (4) is said to have index $1$. If an equation has three radial orbits, with indices $n_{i}$,
$\frac{1}{n_{1}}+\frac{1}{n_{2}}+\frac{1}{n_{3}}=1.$
The classification of the equations (3) with three different radial orbits amounts to solving this equation as a Diophantine one (admitting that $n_{i}$ can take the value $\infty$), from which the last six lines of Table 2 are obtained (the linear equivalence class of such an equation is determined by its indices, and the equations thus obtained happen to be univalent). The other lines of this table follow from organizing quadratic homogeneous vector fields according to the multiplicities of their radial orbits, to the indices of the simple ones, and from imposing the conditions guaranteeing the simplicity of the poles of the associated equation (4).
2.2. Canonical equations and Malmquist’s conditions
Consider the system of equations:
$\begin{split}y\frac{\mathrm{d}x}{\mathrm{d}t}&=na(t)x+b(t)y+\sum_{i+j\geq 2}f_{ij}(t)x^{i}y^{j},\\ \frac{\mathrm{d}y}{\mathrm{d}t}&=a(t)+\sum_{i+j\geq 1}g_{ij}(t)x^{i}y^{j},\end{split}$ (5)
with holomorphic right-hand side, with $n\in\mathbf{Z}$, $n\geq 1$, and with $a$ a non-vanishing function. Following Bureau, we will call these equations canonical. The integer $n$ appearing in such an equation will (also) be called its index. We have the following result, appearing in [Mal21, §11] (see also [Bur64, §2, Thm. IV]), which was probably well-known before Malmquist.
Proposition 6 (Malmquist).
In the case $n=1$, for the system (5) to be free of movable critical points, it is necessary that $b=0$.
We refer the reader to Malmquist’s work for a complete proof. The idea is, first, to establish, via a theorem of Dulac, that for every $t_{0}$ there exists an invariant surface in the extended phase space of the form $t=F(x,y)$ with $F(0,0)=t_{0}$, and that, in restriction to this surface, $x$, as a function of $y$, satisfies an equation of Briot-Bouquet type: $y\,\mathrm{d}x/\mathrm{d}y=x+\cdots$; the solution of the latter will have logarithmic terms unless $b(t_{0})=0$. Since this should happen for all values of $t$, we have that $b$ vanishes identically.
If $n>1$, we may, in the extended phase space, blow up the curve $(x,y)=(0,0)$: for
$u=\frac{b}{(n-1)a},$
and for $z$ defined by $x=(z-u)y$, the system becomes
$\begin{split}y\frac{\mathrm{d}z}{\mathrm{d}t}&=(n-1)a\xi+yh(z,y),\\ \frac{\mathrm{d}y}{\mathrm{d}t}&=a+\sum_{i+j\geq 1}g_{ij}(z-u)^{i}y^{i+j},\end{split}$
for
$h(z,y)=u^{\prime}+\sum_{i+j\geq 2}f_{ij}(z-u)^{i}y^{i+j-2}-\sum_{i+j\geq 1}g_{ij}(z-u)^{i+1}y^{i+j-1}.$
The transformed system is still a canonical one, but, in it, the index $n$ has decreased by one. By iterating this procedure, we arrive to the case $n=1$, where the criterion of Proposition 6 may be applied.
3. The classification
We follow Bureau’s steps for the greater part of the classification. We start with an equation of the form
$\begin{split}y^{\prime}&=P(y,z,t)y+Ay+Bz+a,\\ z^{\prime}&=Q(y,z,t)z+Cy+Dz+b,\end{split}$ (6)
free of movable critical points, with $P$ and $Q$ quadratic homogeneous polynomials in $y$ and $z$ (at least one of them non-trivial), whose coefficients are holomorphic functions of the independent variable, defined in a domain $U\subset\mathbf{C}$. By the results in Section 2.1, for all $t$ in $U$, the linear equivalence class of the quadratic part of (6) is represented by one of the autonomous systems of Table 2, and is thus, with the exception of a discrete set of values of $t$, locally constant (this is the first source of exceptional values of $t$ in Theorem 1). By restricting $t$ to a neighborhood of such a value, up to a change of coordinates that is linear in $y$ and $z$, and that preserves the independent variable, we may suppose that the quadratic part in (6) is independent of $t$, and that it is one of the quadratic systems of Table 2.
By applying suitable transformations that are affine on $y$ and $z$, possibly involving a change of independent variable, we will simplify the system without modifying its quadratic part. We will consider changes of coordinates of the forms
$y$ $=\lambda Y+h,\quad$ $z$ $=\mu Z+k,$ $t$ $=\varphi(\tau);$ ($\text{T}_{1}$)
$y$ $=\lambda Y+h,\quad$ $z$ $=\lambda Z+k,$ $t$ $=\varphi(\tau);$ ($\text{T}_{2}$)
$y$ $=\lambda Y+h,\quad$ $z$ $=\mu Z+\nu Y+k,\quad$ $t$ $=\varphi(\tau),$ ($\text{T}_{3}$)
where the parameters $\lambda$, $\mu$, $\nu$, $h$ and $k$ are functions of $t$, chosen to solve a particular system of algebro-differential equations in a neighborhood of some $t_{0}\in U$ (the need, in Theorem 1, to restrict the domains is also a consequence of the local character of this analysis).
For systems other than system I, when extending $\mathbf{C}^{2}\times U$ into $\mathbf{P}^{2}\times U$ by
$(y,z,t)\mapsto(y:z:1,t),$ (7)
the system will have poles along the line at infinity for all values of $t$, with singularities at the directions of the radial orbits of the quadratic part. At a singularity associated to a radial orbit with positive index $n$, we will obtain a canonical equation (5) with index $n$, and the results in Section 2.2 will impose a necessary condition for the absence of movable critical points, of algebro-differential nature, which may be explicitly calculated after performing $n-1$ blow-ups. We will meet some changes of coordinates that may be framed within this setting; for instance, the coordinates $(1/y,z)$ amount to considering, first, a change of affine chart in $\mathbf{P}^{2}$, $(u,v)=(1/y,z/y)$, followed by a blow-up, $(u,v)\mapsto(u,v/u)=(1/y,z)$.
In some cases, we will, after having calculated all or some of these conditions, relate the system to a second-order equation in one variable, and join the Painlevé-Gambier classification at an advanced stage, either to obtain conditions for the absence of movable critical points, or to integrate the equations. In this we differ from Bureau, who opts for calculating as many necessary conditions as possible directly from the system.
We will thus obtain systems with few parameters, which, in most cases, we will be able either to integrate, or to relate to a time-tested equation already known to be free of movable critical points.
We now begin the classification. Systems III and IV will be studied together with system IX. As discussed in Remark 5, there is no system X.
3.1. System I
Studied by Bureau in [Bur81b, §9]. It represents the systems of the form
$\begin{split}y^{\prime}&=L(y,z,t)y+Ay+Bz+a,\\ z^{\prime}&=L(y,z,t)z+Cy+Dz+b,\end{split}$
with $L$ a linear form in $y$ and $z$. These have long been considered as natural generalizations of the Riccati equation (see [Mal23, §31], [Gou29, Ch. XXI]): when extending the phase space by (7), the system extends holomorphically. Their solutions are thus of the form
$(y(t),z(t))=\left(\frac{\alpha y_{0}+\beta z_{0}+\gamma}{\eta y_{0}+\theta z_{0}+\iota},\frac{\delta y_{0}+\epsilon z_{0}+\zeta}{\eta y_{0}+\theta z_{0}+\iota}\right),$
with $\alpha$, $\beta$, etc. functions of $t$, and these equations are hence naturally free of movable critical points in all the domain where they are defined. A linear transformation redresses $L(y,z,t)$ into $-y$, and, after this, a suitable transformation $\text{T}_{3}$ brings the system into
$\begin{split}y^{\prime}&=-y^{2}+Bz,\\ z^{\prime}&=-yz\end{split}$
(a form simpler than Bureau’s, who uses the more restrictive transformation $\text{T}_{2}$). Explicit expressions for its integration follow from the fact that $(1/z)^{\prime\prime}=B$.
3.2. System II
Studied by Bureau in [Bur81b, §10]. By means of a transformation $\text{T}_{2}$, the system may be cast into Bureau’s form
$\begin{split}y^{\prime}&=y(y-2z)+Ay+a,\\ z^{\prime}&=-z^{2}+Cy.\end{split}$
For $y=1/u$ and $z=v/u$ we have the canonical equation of index $1$
$\begin{split}u^{\prime}&=-1-Au+2v-au^{2},\\ uv^{\prime}&=-v+Cu-Auv+v^{2}-au^{2}v.\end{split}$
From Proposition 6, $C=0$ is a necessary condition for the absence of movable critical points. We will henceforth assume it. At the point $(u,v)=(0,1)$, the system is a canonical equation of index $1$, from which we obtain the necessary condition $A=0$. With it, $u=y-z$ solves the Riccati equation with holomorphic coefficients $u^{\prime}=u^{2}+a$, which is free of movable critical points. Thus, the system solved by $u$ and $z$ (and hence the one solved by $y$ and $z$) is free of movable critical points.
3.3. System V
Studied by Bureau in [Bur81b, §13]. The general form of the system is
$\begin{split}y^{\prime}&=Ay+Bz+a,\\ z^{\prime}&=6y^{2}+Cy+Dz+b.\end{split}$ (8)
Let us not spare details in order to discuss Bureau’s omission in this case. The most general affine transformation preserving the systems (8) is of type $\text{T}_{3}$. By choosing it in the form
$y=\lambda Y+h,\quad z=\mu Z+\lambda\nu Y+k,\quad\frac{\mathrm{d}t}{\mathrm{d}\tau}=\frac{\mu}{\lambda^{2}},$
with $\lambda$, $\mu$, $\nu$, $h$ and $k$ functions of $t$, we obtain
$\frac{\mathrm{d}Y}{\mathrm{d}\tau}=\frac{\mu}{\lambda^{2}}\left(A+\nu B-\frac{\lambda^{\prime}}{\lambda}\right)Y+\frac{\mu^{2}}{\lambda^{3}}BZ+\frac{\mu}{\lambda^{3}}(a+hA+kB-h^{\prime}),\\ \frac{\mathrm{d}Z}{\mathrm{d}\tau}=6Y^{2}+\frac{1}{\lambda}(C-\nu A-\nu^{2}B+\nu D-\nu^{\prime}+12h)Y+\frac{\mu}{\lambda^{2}}\left(D-\nu B-\frac{\mu^{\prime}}{\mu}\right)Z\\ +\frac{1}{\lambda^{2}}(b+hC+kD+6h^{2}-k^{\prime}-\nu[a+hA+kB-h^{\prime}]).$
($f^{\prime}$ denotes the derivative of $f$ with respect to $t$.) This shows that, in (8), the condition $B=0$ is, with respect to affine changes of coordinates, an intrinsic one. We leave out of our analysis the values of $t$ where $B$ has an isolated zero, and restrict to domains where either $B$ vanishes identically, or where it does nor vanish at all (this is another source of special values of the independent variable in the statement of Theorem 1; conditions of the sort will be implicitly assumed throughout the rest of the article).
If $B\neq 0$, the system may be brought into Bureau’s form
$\begin{split}y^{\prime}&=z,\\ z^{\prime}&=6y^{2}+f,\end{split}$ (9)
equivalent to the second-order equation
$y^{\prime\prime}=6y^{2}+f.$ (10)
The latter belongs to the class leading to the first Painlevé equation, studied in detail by Painlevé himself in [Pai00, §15] (see also [Inc44, §14.311, p. 328–329]). For it, Painlevé found that the necessary and sufficient condition for the absence of movable critical points is $f^{\prime\prime}=0$. If $f^{\prime}\neq 0$, the equation is, up to translations in the independent variable and scalings, the equation defining the first Painlevé transcendent $y^{\prime\prime}=6y^{2}+t$. If $f$ is the constant $f_{0}$, Eq. (10) has the first integral $(y^{\prime})^{2}=4y^{3}+2f_{0}y+K$, and is solved by elliptic functions and their rational degenerations. In this case, $4y^{3}+2f_{0}y-z^{2}$ is a first integral for (9); its generic level surfaces are elliptic fibrations over the domain of the independent variable.
Bureau obtains the necessary condition $f^{\prime\prime}=0$ through a canonical equation obtained from Eq. (9). This works both ways: Bureau’s analysis implies that, in Eq. (10), $f^{\prime\prime}=0$ is a necessary condition for the absence of movable critical points. We will reobtain this condition in sections 3.7.2.1, 3.7.2.3 and 3.11.
The case $B=0$ is overlooked by Bureau. In it, in the original system (8), $y$ solves a linear equation with holomorphic coefficients, and so does $z$; the system is free of movable critical points. A suitable transformation $\text{T}_{3}$ reduces the system to system $\text{V}_{0}$ in Table 1.
3.4. System VI
Studied by Bureau in [Bur81b, §14]. It is a limit of system VIII as $n\to\infty$. The system can be brought into Bureau’s form
$\begin{split}y^{\prime}&=Bz+a,\\ z^{\prime}&=z(z+y)+b,\end{split}$
by a transformation of type $\text{T}_{2}$. In the coordinates $u=-1/z$ and $v=y/z$, we find the canonical equation of index $1$,
$\begin{split}u^{\prime}&=1+v+bu^{2},\\ uv^{\prime}&=v+Bu-au^{2}+v^{2}+bu^{2}v,\end{split}$
from which we obtain that $B=0$ is a necessary condition for the absence of movable critical points. The condition is also sufficient: with it, $z$ solves a Riccati equation with holomorphic coefficients.
3.5. System VII
Studied by Bureau in [Bur81b, §16]. It is a limit of system IX as $n\to\infty$. Its general form is
$\begin{split}y^{\prime}&=Ay+Bz+a,\\ z^{\prime}&=yz+Cy+Dz+b.\end{split}$ (11)
We consider changes of coordinates of the form $\text{T}_{1}$, for which the condition $B=0$ is an intrinsic one.
If $B\neq 0$, the system may be brought into Bureau’s form
$\begin{split}y^{\prime}&=z+a,\\ z^{\prime}&=yz+b.\end{split}$
In the coordinates $(u,v)$ defined by $y=-1/u$, $z=(v+\frac{1}{2})/u^{2}$, we find the canonical system of index $2$
$\begin{split}u^{\prime}&=\textstyle\frac{1}{2}+v+au^{2},\\ uv^{\prime}&=v+au^{2}+2v^{2}+bu^{3}+2au^{2}v,\end{split}$
from which we obtain that $a=0$ is a necessary condition for the absence of movable critical points. With it, $y^{\prime\prime}=yy^{\prime}+b$, and thus $y$ solves the Riccati equation with holomorphic coefficients $y^{\prime}=\frac{1}{2}y^{2}+\int b$: the system is free of movable critical points.
If $B=0$, in (11), $y$ solves a linear equation with holomorphic coefficients, and so does $z$ afterwards, and the system is free of movable critical points. The system can be redressed, via a transformation $\text{T}_{1}$, into the system $\text{VII}_{0}$ of Table 1.
3.6. System VIII
Studied by Bureau in [Bur81b, §17]. It has a parameter $n\in\mathbf{Z}$, $n\geq 1$. By means of a transformation $\text{T}_{2}$, it may be brought into Bureau’s form
$\begin{split}y^{\prime}&=y[y-(n+2)z]+Ay+a,\\ z^{\prime}&=-z(ny+z)+b.\end{split}$ (12)
In the coordinates $(u,v)$ defined by $y=1/u$, $z=(v+1)/u$, we have the canonical system of index $1$
$\begin{split}u^{\prime}&=(n+1)-Au+(n+2)v-au^{2},\\ uv^{\prime}&=(n+1)v-Au-Auv+(n+1)v^{2}+(b-a)u^{2}-au^{2}v.\end{split}$
A first necessary condition for the absence of movable critical points is thus $A=0$, which we assume from now on. For $y=-1/u$, we have the canonical equation of index $n$
$\begin{split}u^{\prime}&=1+au^{2}+(n+2)uz,\\ uz^{\prime}&=nz+bu-uz^{2},\end{split}$ (13)
and, through it, we find a necessary condition for each value of $n$. For the smallest ones, these are
for $\text{for }n=1,\quad$ $b=0;$
for $\text{for }n=2,\quad$ $b^{\prime}=0;$
for $\text{for }n=3,$ $b^{\prime\prime}+ab-3b^{2}=0;$
for $\text{for }n=4,$ $b^{\prime\prime\prime}+4(a-4b)b^{\prime}+2ba^{\prime}=0;$
it seems difficult to give an explicit form of this condition for all values of $n$.
In order to integrate the system (12) with $A=0$, Bureau transforms it into a particular case of the second-order equation in one variable studied by Gambier in [Gam10, Ch. VII, Eq. 5, p. 53], which he revisits in [Bur64, §49] (when $a\neq 0$, he passes through the second-order equation satisfied by $1/y$; when $n\neq 2$ and $b\neq 0$, through the one satisfied by $z$; he deals separately with the case $n=2$). Gambier’s equation has a rather convoluted scheme of integration. In it, the necessary and sufficient conditions for the absence of movable critical points are easy to establish for the equations corresponding to low values of $n$; for higher values, either the absence of movable critical points can be established directly, through the vanishing of a particular explicit expression, or the equation may be transformed into one of the same family, in which $n$ is replaced by $n-2$. We refer to Gambier’s work for details.
The eventual movable critical points of the system are easy to locate. For $z$ and $w=y-z$, we have the system
$\begin{split}w^{\prime}&=w^{2}+a-b,\\ z^{\prime}&=-(n+1)z^{2}-nwz+b.\end{split}$
In it, $w$ solves a Riccati equation with holomorphic coefficients, and has thus meromorphic solutions, with movable poles; $z$ solves a Riccati equation with meromorphic coefficients, in which the movable critical points, if any, come from the movable poles of $w$.
With the condition $A=0$, the necessary condition for the absence of movable critical points coming from the canonical equation (13) turns out to be a sufficient one. We will establish this geometrically, without integrating the equations, in connection with the ideas in [GR97, Sect. 3.4]. Embed $\mathbf{C}^{2}$ into $\mathbf{P}^{1}\times\mathbf{P}^{1}$ via $(w,z)\to(w:1,z:1)$. Let $u=1/w$ and $v=1/z$, so that $(w:1,1:v)$, $(1:u,z:1)$ and $(1:u,1:v)$ give the three other charts of $\mathbf{P}^{1}\times\mathbf{P}^{1}$. In these,
$\begin{split}w^{\prime}&=w^{2}+a-b,\\ v^{\prime}&=n+1+nvw-bv^{2};\end{split}$
$\begin{split}u^{\prime}&=-1+(b-a)u^{2},\\ uz^{\prime}&=-nz+bu-(n+1)uz^{2};\end{split}$ (14)
$\begin{split}u^{\prime}&=-1+(b-a)u^{2},\\ uv^{\prime}&=nv+(n+1)u-buv^{2}.\end{split}$ (15)
The system has poles along $u=0$, and nowhere else. In the chart (14), the equation has a singularity with index $n$ along $p:(u,z)=(0,0)$, and, in the chart (15), a singularity with index $-n$ at $q:(u,v)=(0,0)$.
Elementary transformations play an important role in the birational theory both of ruled surfaces and of the Riccati equation (see [BHPV04, Ch. 6, Sect. 7] and [Bru15, Ch. 4, §1]; elementary transformations are called “flipping the fiber” in the latter). We will perform parametric (non-autonomous) versions of such transformations.
Start with the rational fibration $\Pi:\mathbf{P}^{1}\times\mathbf{P}^{1}\times U\to\mathbf{P}^{1}\times U$, $(1:u,z:1,t)\mapsto(1:u,t)$. Let $F$ be the fiber of $\Pi$ above $\{u=0\}\times U$. For the intersection with a slice $\Sigma:t=t_{0}$, $F\cap\Sigma$ is, within $\Sigma$, a rational curve of vanishing self-intersection. The blowing up of $p$ produces a surface $E$, and transforms $F$ into a surface $F^{\prime}$. Within the transform of $\Sigma$, both $E\cap\Sigma$ and $F^{\prime}\cap\Sigma$ are rational curves of self-intersection $-1$. By the parametric version of Grauert’s criterion due to Nakano and Fujiki [Nak71, FN72], $F^{\prime}$ may be contracted along the fibers of $t$, producing a curve $q^{\prime}$ transverse to $t$. This transforms $E$ into a surface $E^{\prime}$, and, within the new transform of $\Sigma$, $E^{\prime}\cap\Sigma$ is a curve of vanishing self-intersection. The total space $\mathbf{P}^{1}\times\mathbf{P}^{1}\times U$ has been modified into a threefold $M$, and the transform of $\Pi$ subsists as a locally trivial rational bundle $\Pi^{\prime}:M\to\mathbf{P}^{1}\times U$. The fiber $F$ has been replaced by $E^{\prime}$, but all the other fibers have gone by unchanged. The transformed equation has poles along $E^{\prime}$. There is a curve $p^{\prime}$ on $E^{\prime}$, along which we have a canonical equation with index $n-1$; at the curve $q^{\prime}$, we find a canonical equation with index $1-n$. This procedure can be iterated: when applied to a point of index $1$, the conditions of Proposition 6 are equivalent to the fact that the poles in the equation disappear after blowing up the point, thus removing all poles from the equation.
After iterating the procedure, $\mathbf{P}^{1}\times\mathbf{P}^{1}\times U$ has been birationally transformed into a threefold where the equation extends holomorphically (where it no longer has poles), such that the fibers of the projection onto $t$ compactify the original $\mathbf{C}^{2}\times\{t\}$. This implies that the equation is free of movable critical points (see [Oka79, §1], where Okamoto resorts to Ehresmann’s theorem, or Painlevé’s 17th of his Leçons de Stockholm [Pai97]).
For example, when $b=0$, starting from the chart $(1:u,z:1)$, after the composition of $n$ elementary transformations given by $z=su^{n}$, the system transforms into
$\begin{split}u^{\prime}&=-1-au^{2},\\ s^{\prime}&=nasu-(n+1)u^{n}s^{2},\end{split}$
one which no longer has poles. (In particular, for all values of $n$, the condition $b=0$ guarantees the absence of movable critical points.)
3.7. System IX (plus systems III and IV)
These were studied by Bureau in §11, §12 and §18 of [Bur81b]. Their general form is
$\begin{split}y^{\prime}&=-y^{2}+Ay+Bz+a,\\ z^{\prime}&=(n-1)yz+Cy+Dz+b,\end{split}$ (16)
with $n\in\mathbf{Z}$, $n\neq 0$. System III corresponds to $n=-1$, system IV to $n=1$, and system X to the remaining cases. The latter is split in two: system IX.A corresponds to $n<-1$, and system IX.B to $n>1$. Warning: in system IX.A, Bureau’s normalization differs from the one here used by the sign of $n$.
Let us be overly explicit in the calculations in this part, in which Bureau’s analysis presents some oversights. Under a transformation of type $\text{T}_{1}$, in the form
$y=\lambda Y+h,\quad z=\mu Z+k,\quad\frac{\mathrm{d}t}{\mathrm{d}\tau}=\frac{1}{\lambda},$
with $\lambda$, $\mu$, $h$ and $k$ functions of $t$, we have
$\frac{\mathrm{d}Y}{\mathrm{d}\tau}=-Y^{2}+\frac{1}{\lambda}\left(A-\frac{\lambda^{\prime}}{\lambda}-2h\right)Y+\frac{\mu}{\lambda^{2}}BZ+\frac{1}{\lambda^{2}}(a+hA+kB-h^{2}-h^{\prime}),\\ \frac{\mathrm{d}Z}{\mathrm{d}\tau}=(n-1)YZ+\frac{1}{\mu}(C+(n-1)k)Y+\frac{1}{\lambda}\left(D-\frac{\mu^{\prime}}{\mu}+(n-1)h\right)Z+\\ +\frac{1}{\lambda\mu}(b+(n-1)hk+hC+kD-k^{\prime})$
($f^{\prime}$ denotes the derivative of $f$ with respect to $t$).
Remark 7.
Under such changes of coordinates, we have that, in (16),
  • the condition $B=0$ is an intrinsic one; and that
  • if $n=-3$ and $B\neq 0$, so is the condition $02A-D-B^{\prime}/B=$ 0.
Our analysis will be split into the following cases:
3.7.1. Case $n=1$ (system IV)
The system can be brought, by a transformation $\text{T}_{1}$ as above, into the form
$\begin{split}y^{\prime}&=-y^{2}+Bz,\\ z^{\prime}&=Cy\end{split}$ (17)
(a form simpler than Bureau’s, who uses the more restrictive transformation $\text{T}_{2}$). In the coordinates $u=1/y$, $v=z/y$, it reads
$\begin{split}u^{\prime}&=1-Bvu,\\ uv^{\prime}&=v+Cu-Buv^{2},\end{split}$
a canonical equation of index $1$, from which we obtain the necessary condition $C=0$. With it, system (17) reduces to a Riccati equation in $y$ with holomorphic coefficients, and is thus free of movable critical points.
We will henceforth suppose that $n\neq 1$.
3.7.2. Case $B\neq 0$, $n\neq-3$
Up to a suitable transformation as above, we may suppose that the system has Bureau’s form
$\begin{split}y^{\prime}&=-y^{2}+z+a,\\ z^{\prime}&=(n-1)yz+b.\end{split}$ (18)
When $y$ and $z$ are respectively given the weights $1$ and $2$, its quasihomogeneous component of highest degree is
$\begin{split}y^{\prime}&=-y^{2}+z,\\ z^{\prime}&=(n-1)yz.\end{split}$ (19)
In this way, (18) can be seen as a lower-order perturbation of (19), and the whole procedure described for quadratic equations can be adapted to the study of such systems. For instance, if (18) is free of movable critical points, (19) has only single-valued solutions. For $v=z/y^{2}$, we have, from (19), the Briot-Bouquet equation
$-\frac{v^{\prime\prime}}{(v^{\prime})^{2}}=-\frac{3v-n}{v(2v-n-1)},$ (20)
discussed in Section 2.1. We must impose the condition $n\neq-1$ in order for the pole at $v=0$ to be simple. The expression (20) has a simple pole at
$v_{1}=\frac{1}{2}(n+1),$
with residue
$\frac{n-1}{2(n+1)}-1.$
If $v$ is univalent, the latter is either equal to $-1$ or of the form $11/k-$ 1 for $k\in\mathbf{Z}$, this is, $2(n+1)/(n-1)=2+4/(n-1)$ is an integer (the index). The admissible cases are $n\in\{2,3,5\}$. The indices of the poles of (20) appear in Table 3.
$n$ 0 $v_{1}$infty
$2$ $3$ $6$ $2$
$3$ $4$ $4$ $2$
$5$ $6$ $3$ $2$
Table 3. Indices of the Briot-Bouquet equation associated to equation (19).
Considering, with Bureau, the coordinates $u=1/y$, $v=z/y^{2}$, we have
$\begin{split}u^{\prime}&=1-v-au^{2},\\ uv^{\prime}&=(n+1)v-2v^{2}+bu^{3}-2au^{2}b.\end{split}$ (21)
We have two points along $u=0$: $v=0$ and $v=v_{1}$. At each one of them we have a canonical equation whose index appears in Table 3. This will give, in each of the admissible cases, two necessary conditions for the absence of movable critical points.
3.7.2.1 Subcase $n=2$
The necessary condition coming from (21) along $(u,v)=(0,0)$ is $b=0$, which we assume from now on. Set $a=12q$. For $u=\frac{1}{6}z+q$ and $f=q^{\prime\prime}-6q^{2}$, we have that $u^{\prime\prime}=6u^{2}+f$, the equation appearing in Section 3.3 as Eq. (10), related to the first Painlevé equation. As we there discussed, the necessary and sufficient condition for the absence of movable critical points is $f^{\prime\prime}=0$; this is also the necessary condition arising from Eq. (21) at $(u,v)=(0,v_{1})$.
The equation is birationally equivalent to the natural first-order system in one variable associated to the first Painlevé equation: $u$ and $u^{\prime}$ may be rationally expressed in terms of $y$ and $z$; on their turn, from (18), $y$ and $z$ may be rationally expressed in terms of $u$ and $u^{\prime}$.
3.7.2.2 Subcase $n=3$
We have, for $f=-2a$ and $g=b+a^{\prime}$,
$y^{\prime\prime}=2y^{3}+fy+g.$ (22)
We find in [Inc44, §14.315, p. 333] that the conditions for this equation to be free of movable critical points are $f^{\prime\prime}=0$ and $g^{\prime}=0$ ($a^{\prime\prime}=0$ and $b^{\prime}=0$). These are exactly the two conditions arising from equation (21) (they will be also discussed in Section 3.10). Under these, (22) is the general form of the equation leading to the Painlevé II transcendent [Pai02, §11]. If $f^{\prime}\neq 0$, up to a translation the independent variable and suitable scalings, equation (22) gives one of the standard forms of the Painlevé II equation
$y^{\prime\prime}=2y^{3}+ty+g_{0}.$
When $f^{\prime}=0$, (22) is integrated by elliptic functions and their rational degenerations: if $f$ and $g$ are, respectively, the constants $f_{0}$ and $g_{0}$, we have the first integral $(y^{\prime})^{2}=y^{4}+f_{0}y^{2}+2g_{0}y+K$ (see also [Inc44, §14.315, p. 333]). When substituting for $y^{\prime}$ from (18), we obtain that
$2y^{2}z-z^{2}+2by-2az-a^{2}$
is a first integral for (18). Its generic level surfaces are elliptic fibrations over the domain of the independent variable.
The system is polynomially equivalent to the first-order system in two variables associated to the general Painlevé II equation: from (18), $y$ and $z$ may be polynomially expressed in terms of $y$ and $y^{\prime}$, and vice versa.
3.7.2.3 Subcase $n=5$
The necessary condition arising from Eq. (21) at $(u,v)=(0,v_{1})$ is $b=-3a^{\prime}$. Set $a=3q$ and $b=-9q^{\prime}$. For $u=-2y$, we have
$u^{\prime\prime}=-uu^{\prime}+u^{3}-12qu+12q^{\prime}.$ (23)
This equation is integrated in [Pai02, p. 23, Eq. 6]: if $w$ is a solution to
$w^{\prime\prime}-6w^{2}=q^{\prime\prime}-6q^{2}$ (24)
different from $q$,
$u=\frac{w^{\prime}-q^{\prime}}{w-q}$ (25)
is a solution to (23). When considered as first-order equations in two variables, equations (23) and (24) are birationally equivalent, since, from (25), $w$ and $w^{\prime}$ can be expressed rationally in terms of $u$ and $u^{\prime}$:
$\begin{split}w&=\textstyle{\frac{1}{6}}u^{2}-q+\textstyle{\frac{1}{6}}u^{\prime},\\ w^{\prime}&=\textstyle{\frac{1}{6}}u^{3}-2uq+\textstyle{\frac{1}{6}}uu^{\prime}+q^{\prime},\end{split}$
and vice versa.
As discussed in Section 3.3, the necessary and sufficient condition for (24) to be free of movable critical points is
$(q^{\prime\prime}-6q^{2})^{\prime\prime}=0,$ (26)
in which (24) becomes the first Painlevé equation or one of its degenerations. Thus, condition (26), is a necessary and sufficient condition for (23) to be free of movable critical points as well (compare with [Inc44, §14.314, p. 331–333]). Condition (26) is also the one arising from the canonical equation (21) at $(u,v)=(0,0)$.
3.7.3. Case $B\neq 0$, $n=-3$
In this case, owing to the condition given in the second item of Remark 7, the system (16) cannot always, despite Bureau’s claim, be brought into the form (18). It can nevertheless be put into the form
$\begin{split}y^{\prime}&=-y^{2}+z,\\ z^{\prime}&=-4yz+Dz+b.\end{split}$
For $u=1/y$, $v=z/y^{2}+1$,
$\begin{split}u^{\prime}&=2-w,\\ uw^{\prime}&=2v-Du+Duv-2v^{2}+bu^{3},\end{split}$
and we thus find the necessary condition $D=0$ (Bureau’s analysis of this case fails to see that there are cases which are not free of movable critical points). We have
$y^{\prime\prime}+6yy^{\prime}+4y^{3}-b=0.$
For $y=\frac{1}{2}w^{\prime}/w$, this equation reduces to the third-order linear equation $w^{\prime\prime\prime}=2bw$, and is thus free of movable critical points.
3.7.4. Case $B=0$
This case was overlooked by Bureau. Through a transformation $\text{T}_{1}$, the system may be brought into the form
$\begin{split}y^{\prime}&=-y^{2},\\ z^{\prime}&=(n-1)yz+Cy.\end{split}$
After solving for $y$, the system reduces to a linear inhomogeneous equation in $z$ with meromorphic coefficients. Its solutions may have movable critical points, arising from the poles of $y$. Its non-constant solutions are $y(t)=1/(t-t_{0})$,
$z(t)=(t-t_{0})^{n-1}\int^{t}\frac{C(s)\,\mathrm{d}s}{(s-t_{0})^{n}}.$
[…]
System III corresponds to the case $n=-1$. In it, by using a transformation $\text{T}_{1}$ instead of one of type $\text{T}_{2}$, we have obtained a normal form simpler than Bureau’s. The one appearing in Table 1 is obtained by changing the sign of $y$.
3.8. System XI
Studied by Bureau in [Bur81b, §21]. A transformation $\text{T}_{2}$ brings the system into the form
$\begin{split}y^{\prime}&=y(y-z)+qy+a,\\ z^{\prime}&=z(z-y)-qz+b.\end{split}$ (27)
(This normal form is simpler than Bureau’s, and exhibits the natural symmetries of the system.) For $u=-1/y$, we have the canonical equation of index $1$
$\begin{split}u^{\prime}&=1-qu+uz+au^{2},\\ uz^{\prime}&=z+bu-quz+uz^{2},\end{split}$
from which we obtain the necessary condition $b=0$. By the symmetry of the system (27) consisting in exchanging $y$ and $z$, $a=0$ is a necessary condition as well. With these conditions,
$y^{\prime\prime}=2yy^{\prime}+qy^{\prime}+q^{\prime}y.$
The latter has the first integral $y^{\prime}=y^{2}+qy+K$, and thus $y$ solves a Riccati equation with holomorphic coefficients; the system is free of movable critical points.
3.9. System XII
Studied by Bureau in [Bur81b, §22]. By means of a transformation $\text{T}_{2}$, the system may be brought into the form
$\begin{split}y^{\prime}&=y(2z+y)+2fy-a,\\ z^{\prime}&=-z(2y+z)-2fz+b.\end{split}$ (28)
For $u=-1/y$, we have
$\begin{split}u^{\prime}&=1-2fu-2uz-au^{2},\\ uz^{\prime}&=2z+bu-2fuz-uz^{2},\end{split}$
a canonical equation of index $2$, from which a necessary condition for the absence of movable critical points is found to be $b^{\prime}=0$. By the symmetries of (28), more evident when $z$ is replaced by $-z$, we also have the condition $a^{\prime}=0$. The involution $(y,z)\mapsto(-y,z+y+2f)$ preserves the general form of the systems (28), and acts upon their coefficients by
$(f,a,b)\mapsto(-f,-a,2f^{\prime}-a+b);$
in this way, we find that $f^{\prime\prime}=0$ is a third necessary condition for the absence of movable critical points. With these three, $y$ is a solution to Eq. (2) for the same $f$, for $\alpha=\frac{1}{2}a-b-f^{\prime}$ and $\beta=a$. Equation (2) is free of movable critical points: it is the general form of the equation leading to the Painlevé IV equation [Gam10, Ch. 1, §9]. When $f^{\prime}\neq 0$, up to translations in the independent variable and scalings, we may suppose that $f(t)=t$, recovering the Painlevé IV equation in its full generality. The cases where $f^{\prime}=0$ are integrated by elliptic functions: if $f$ is the constant $f_{0}$, Eq. (2) has the first integral $K$ defined by
$(y^{\prime})^{2}=y^{4}+4f_{0}y^{3}+4(f_{0}^{2}-\alpha)y^{2}+Ky+\beta^{2}$
(see also [Inc44, §14.331, case 6, p. 339]). In this case, (28) has the first integral
$y^{2}z+yz^{2}+2f_{0}yz-by-az-af_{0},$
whose generic fibers are elliptic fibrations over the domain of the independent variable.
System (28), with the integrability conditions, is birationally equivalent to the first-order rational system in two variables associated to Eq. (2), for, on the one hand, $y^{\prime}$ is a polynomial in $y$ and $z$, and, on the other, through (28), $z$ can be rationally expressed in terms of $y$ and $y^{\prime}$.
3.10. System XIII
Studied by Bureau in [Bur81c, §23–24]. Up to a transformation $\text{T}_{2}$, we may suppose that the system has the form
$\begin{split}y^{\prime}&=\textstyle{\frac{1}{2}}y(2z-y)+2py+a,\\ z^{\prime}&=\textstyle{\frac{1}{2}}z(3y-2z)-4pz+b.\end{split}$ (29)
(It is different from Bureau’s.) For $u=1/z$, we have the canonical equation of index $1$
$\begin{split}u^{\prime}&=\textstyle 1+4pu-bu^{2}-\frac{3}{2}uy,\\ uy^{\prime}&=\textstyle y+au+2puy-\frac{1}{2}uy^{2}.\end{split}$
Hence, $a=0$ is a necessary condition for the absence of movable critical points. With it, $u=\frac{1}{2}y-p$ solves Eq. (22) for
$\begin{split}f&=2p^{\prime}-2p^{2}+b,\\ g&=-p^{\prime\prime}+2pp^{\prime}+bp.\end{split}$ (30)
Following the discussion in Section 3.7.2.2, the necessary and sufficient conditions for Eq. (22) to be free of movable critical points are $f^{\prime\prime}=0$ and $g^{\prime}=0$, in which case we obtain the Painlevé II equation and some of its degenerations. Since, from (30),
$\begin{split}b&=f-2p^{\prime}+2p^{2},\\ p^{\prime\prime}&=2p^{3}+fp-g,\end{split}$
we have that, with the above conditions, $p$ is a solution to another instance of Eq. (22), one with the same $f$ and an opposite sign for $g$.
System (29), with the above integrability conditions, is birationally equivalent to the first-order system in two variables associated to Eq. (22), for $u$ and $u^{\prime}$ can be rationally expressed in terms of $y$ and $z$, and vice versa.
The conditions $f^{\prime\prime}=0$ and $g^{\prime}=0$ can be obtained through some canonical equations obtained from system (29) with $a=0$. For $u=2/y$, we have the canonical equation with index $3$
$\begin{split}u^{\prime}&=1-2up-uz,\\ uz^{\prime}&=3z+bu-uz^{2}-4puz,\end{split}$
from which we obtain the necessary condition for the absence of movable critical points $f^{\prime\prime}+2g^{\prime}=0$. The systems (29) with $a=0$ are preserved by the involution $(y,z)\mapsto(-y,z-y+4p)$, which acts upon their coefficients by $(p,b)\mapsto(-p,b+4p^{\prime})$. The involution acts upon $f$ and $g$ by $(f,g)\mapsto(f,-g)$. Hence, we also have the condition $f^{\prime\prime}-2g^{\prime}=0$. Together, these conditions give the previously discussed ones $f^{\prime\prime}=0$ and $g^{\prime}=0$.
3.11. System XIV
Studied by Bureau in [Bur81c, §25–27]. A transformation $\text{T}_{2}$ brings the system to the form
$\begin{split}y^{\prime}&=y(2z-y)+3py+a,\\ z^{\prime}&=z(y-z)-2pz+b\end{split}$ (31)
(different from Bureau’s). For $(u,v)=(1/y,z/y)$, we have the canonical equation of index $2$
$\begin{split}u^{\prime}&=1-3pu-2v-au^{2},\\ uz^{\prime}&=2v+bu^{2}-5puv-3v^{2}-au^{2}v,\end{split}$ (32)
from which we obtain the necessary condition $b=0$. With it, for $u=1/z$, we have the canonical system of index $2$
$\begin{split}u^{\prime}&=1+2pu-uv,\\ uv^{\prime}&=2v+au+3puv-uv^{2},\end{split}$
and, from it, a second necessary condition, $a^{\prime}=pa$, follows. Assuming these two, we find that $u=z+p$ is a solution to (23) for $q=\frac{1}{12}(p^{\prime}+p^{2}-a)$. By eliminating $a$ from $q$ and $q^{\prime}$, we have that $p$ is a particular solution to (23). System (31) and the first-order system associated to (23) are birationally equivalent. As discussed in Section 3.7.2.3, the necessary and sufficient condition for Eq. (23) to be free of movable critical points is $(q^{\prime\prime}-6q^{2})^{\prime\prime}=0$. Hence, for system (31), assuming the first two necessary conditions, this is a necessary and sufficient condition as well. System V, the one giving the first Painlevé equation, and system XIV are birationally equivalent.
At $(u,v)=(0,2/3)$, Eq. (32) is a canonical system of index $6$, and, assuming the first two necessary conditions for the absence of movable critical points, the condition arising from this point is $(q^{\prime\prime}-6q^{2})^{\prime\prime}=0$; we may thus obtain the necessity of this condition directly from system (31).